OpenStudy (kittiwitti1):

Linear Approximations and Differentials a) Use differentials to find a formula for the approximate volume of a thin cylindrical shell with height h, inner radius r, and thickness Δr. b) What is the error involved in using the formula from part (a)?

OpenStudy (kittiwitti1):



OpenStudy (sooobored):

considering cylindrical coordinates and just looking at the r,theta plane since height/z integral is as simple as multiplying by a constant

OpenStudy (kittiwitti1):

This is Single Variable Calculus.

OpenStudy (sooobored):

 let us assume that dr is a very very small number

OpenStudy (sooobored):

er, i realized, i know how to find it through integration and not differential...

OpenStudy (kittiwitti1):

lol. It's ok, because you tried :-D

OpenStudy (sooobored):

should i continue?

OpenStudy (kittiwitti1):

I have to do it with differentials. sorry :-(

OpenStudy (kittiwitti1):

but thank you for the alternate approach, I will keep it in mind for when I start Calc II :-)

OpenStudy (sooobored):

wait, differential is just the delta r thing right?

OpenStudy (kittiwitti1):

eh... it's something like f prime x I think?\[f'(x)\]

OpenStudy (kittiwitti1):

I'm not sure how to define it 100% but I know generally how to apply differentials I guess

OpenStudy (sooobored):

cor convenience sake\[\Delta r = dr\] ok non calculus approach you obviously know that the the area of a circle is \[A=\pi R^2\] and so going back to the picture, the big circle has a radius of r and the little circle has a radius of r-dr

OpenStudy (sooobored):

since you want to find the small area inbetween the two circles you just subtract areas now you make the assumption that dr is infinitesmally small so we also make the assumption that dr^2 ~0

OpenStudy (sooobored):

then multiply by height to get volume of the cylindrical shell

OpenStudy (kittiwitti1):

Er, Δr is equal to dr?

OpenStudy (sooobored):

well, not easy to type a triangle inline

OpenStudy (kittiwitti1):

oh, I copypaste the delta symbol from the nets. Δ but you can also use \Delta which shows up as \(\Delta\)

OpenStudy (sooobored):

well im pretty inexperienced with the wonders that is latex

OpenStudy (kittiwitti1):

lol I abuse Google in terms of LaTeX :-P

OpenStudy (dumbcow):

to set up the integral use fact that volume is sum of areas (circumference*height) \[V = \int\limits_a^b (2 \pi r) (h) dr\] \[V = \int\limits_R^{R+dr} (2\pi r)(h) dr\]

OpenStudy (sooobored):

oh yea, you could put the differential on the outside the result should still be teh same

OpenStudy (kittiwitti1):

@dumbcow I'm in Single Variable Calculus...

OpenStudy (kittiwitti1):

...I don't know anything about integrals. Zilch. :-(